# Latent Diffusion Prior Enhanced Deep Unfolding for Snapshot Spectral Compressive Imaging

### v1

@@ Abstract
Snapshot compressive spectral imaging reconstruction aims to reconstruct three-dimensional spatial-spectral images from a single-shot two-dimensional compressed measurement. Existing state-of-the-art methods are mostly based on deep unfolding structures but have intrinsic performance bottlenecks: $i$) the ill-posed problem of dealing with heavily degraded measurement, and $ii$) the regression loss-based reconstruction models being prone to recover images with few details. In this paper, we introduce a generative model, namely the latent diffusion model (LDM), to generate degradation-free prior to enhance the regression-based deep unfolding method. Furthermore, to overcome the large computational cost challenge in LDM, we propose a lightweight model to generate knowledge priors in deep unfolding denoiser, and integrate these priors to guide the reconstruction process for compensating high-quality spectral signal details. Numeric and visual comparisons on synthetic and real-world datasets illustrate the superiority of our proposed method in both reconstruction quality and computational efficiency. Code will be released.
@@ 1 Introduction
In contrast to normal RGB images which only have three spectral bands, hyperspectral images (HSIs) contain multiple spectral bands with more diverse spectral information. The spectral information serves to characterize distinct objects assisting high-level image tasks [51, 50, 29, 27, 42] and the observation of the world like medical imaging [32, 49] and remote sensing [15, 33]. However, the capture of HSIs is a question that has been studied for a long time because we need to collect HSI signals by 2D sensors.
@@
Figure 1: (a) Comparison of PSNR (dB)-FLOPs (G) with previous HSI reconstruction methods. Our proposed method outperforms previous methods using even less computational costs. (b) The ablation study of using different time steps in diffusion. Our method can achieve the desired results requires only very few steps.
@@
The conventional way of spectral imaging is scanning. It scans the scenes along one dimension such as the spectral or spatial dimension, consuming plenty of time. This way of capture limits the imaging objects to static ones. Thus, for many years, scientists have focused on how to collect HSIs in a quick and convenient method. In 2007, based on compressive sensing theory, a single-shot compressive spectral imaging way  [14] was created to efficiently collect HSIs, named coded aperture snapshot spectral imaging (CASSI). The later improvement works [53, 38] provide better imaging quality and lower cost. CASSI modulates the HSI signal across various bands and combine all the modulated spectra to produce a 2D compressed measurement. Consequently, the task of reconstructing the 3D HSI signals from the 2D compressive measurements presents a fundamental challenge for the CASSI system.
@@
The reconstruction process can be viewed as solving an ill-posed problem. Many attempts at solving this problem including traditional model-based methods [2, 1, 60] and the learning-based methods[8, 40, 37] have been proposed since the inception of CASSI system. The deep unfolding network is a combination of convex optimization and neural network prior (denoiser), enjoying both the interpretability of the model-based method and the power of learning-based methods. This branch of methods leads the development trend in recent years [56, 6, 39, 12, 28] and achieves state-of-the-art performance.
@@
However, unlike denoising or reconstruction that recovers from natural images, CASSI reconstruction has to recover HSIs from the compressed domain measurements, which results in severe degradation according to physical modulation, spectral compression, and unpredictable system noise. Thus, the CASSI reconstruction problem is much harder to learn intrinsic HSI properties than the normal image restoration tasks. In the unfolding framework of the CASSI reconstruction method, the denoising network plays a critical role in deciding the final performance, which is embedded in each stage of the deep unfolding network. However, it always suffers from the performance bottleneck due to the intrinsic ill-posed problem of dealing with heavily degraded measurements. Therefore, a high-performance denoiser with degradation-free knowledge is desired for CASSI reconstruction. Another problem is that previous popular regression-based reconstruction methods have difficulty in recovering details, because the widely used regression losses are conservative with high-frequency details [46].
@@
To address these challenges, we introduce a generative prior in this paper to guide the reconstruction process in an unfolding framework. During training, the prior will be first learned from clean HSIs by an image encoder and then generated by a Latent Diffusion Model (LDM) from Gaussian noise and compressed measurement. Then, the learned prior is embedded into the deep denoiser of the unfolding network by a prior-guided Transformer. Significantly, our unfolding network is able to leverage external prior knowledge from clean HSIs and the powerful generative ability of LDM enhancing its reconstruction performance. The primary contributions presented in this paper can be summarized as follows:
@@
    We propose a novel LDM-based unfolding network for CASSI reconstruction, where the clean image priors are generated by a latent diffusion model to facilitate high-quality hyperspectral reconstruction. To the best of our knowledge, this is the first attempt to combine the physics-driven deep unfolding with generative LDM in CASSI reconstruction.
@@
    We design a three-in-one Transformer structure dubbed Trident Transformer (TT) to efficiently extract the correlation among prior knowledge, spatial, and spectral for CASSI reconstruction.
@@
Figure 2: The upper row: the error maps of the previous SOAT and our method. Our method exhibits fewer errors in the edges and textures. The lower row: the feature map (drawn by heatmaps) changes before and after applying LDM enhancement. The enhanced features demonstrate increased concentration on significant parts and edges.
@@ 2.1 Diffusion Model in Low-level Vision
Diffusion models (DMs) [18, 47] are probabilistic generative models, which model the data distribution by learning a gradual iterative denoising process from the Gaussian distribution to the data distribution. Notably, they demonstrate promising capabilities in generating high-quality samples that encompass a wide range of modes, including super-resolution [13] and inpainting [34]. In light of the impressive achievements of diffusion models in image domains, numerous research endeavors [16, 19, 3, 20] have extended it to video generation. However, diffusion models suffer from significant computation inefficiency regarding data sampling, primarily due to the iterative denoising process required for inference. To address this challenge, several methods propose effective sampling techniques from trained diffusion models [48, 63], or alternatively learning the data distribution from a low-dimensional latent space [44], i.e. the latent diffusion model. The latent diffusion has a relatively faster speed and powerful generative ability for super-resolution and inpainting, but similar to the normal diffusion model, it is also prone to issues such as misaligned distribution of fine details and the occurrence of unwanted artifacts, leading to suboptimal performance in distortion-based metrics, e.g., PSNR. Moreover, latent diffusion costs large computational resources both for training and inference due to its large-size encoder and denoiser. Towards this end, some works combine the generative diffusion model with the regression restoration network and work well on distortion-based metrics like deblurring [43]. The recent works [9, 58] employ LDM on many low-level vision tasks and achieve state-of-the-art performance with reasonable computational cost. We name these methods ‘integrated diffusion’ to distinguish them from the pure diffusion method. However, it is challenging to efficiently employ diffusion models for compressive hyperspectral image reconstruction with heavily degraded measurements.
@@ 2.2 Hyperspectral Image Reconstruction
Before the advent of the deep learning wave, traditional model-based methods iteratively solved this inverse problem by convex optimization [54, 64, 52] with some hand-crafted constraints based on image priors, like sparsity [26] and low-rank [31]. These methods are robust and interpretable but require manual parameter tuning with low reconstruction speed and performance. With the help of deep learning, Plug-and-play (PnP) algorithms [7, 45, 61, 62], embeds pre-trained denoising networks into convex optimization to solve the reconstruction problem, but still has limitations on performance because of the pre-trained denoiser. In recent years, the End-to-end (E2E) reconstruction directly trains a powerful deep neural network, like convolutional neural network (CNN) [10, 33, 22] and Transformers [5, 4], to learn the recovery process from inputs (measurements) to outputs (desired HSIs). However, this simple design lacks interpretability and robustness for various hardware systems. Therefore, an interpretable design of a reconstruction network that unfolds a convex optimization process named deep unfolding is proposed to leverage these problems. A series of CASSI reconstruction works based on deep unfolding [36, 35, 6, 12, 59, 28] are proposed and become the state-of-the-art method. Deep unfolding is able to combine both interpretability in model-based methods and performance in deep learning-based methods to reconstruct CASSI at a fast speed. It changes iterative steps in optimization into several stages in a single network. The prior for optimization becomes a deep neural network denoiser. Since the unfolding needs to define the forward model of imaging, it is also considered as a physics-driven network. However, the recent unfolding networks still have bottlenecks for their regression-based denoiser design and the difficulty of dealing with compressive measurement features. Bearing these concerns, in this work, we propose an ‘integrated diffusion’ module and integrate it into the physics-driven deep unfolding framework and design an efficient way to aggregate complex features during reconstruction.
@@
Figure 3: (a) The single disperser CASSI imaging process. HSI data cube is captured by a monochromatic sensor. (b) GC-GAP projection. (c) Latent encoder. (d) Our deep unfolding framework of N stages with the priors generated by diffusion model.
@@ 3 Proposed Model
3.1 Problem Formulation
The CASSI system has high efficiency in capturing 3D spectral signals by initially coding spectral data with different wavelengths in an aperture and then integrating them into a 2D monochromatic sensor. The mathematical forward process of the widely used single-disperser CASSI (SD-CASSI) [52] can be illustrated as Fig. 3 (a). As can be seen, the original HSI data, denoted as $\bm{X}\in\mathbb{R}^{W\times H\times N_{\lambda}}$, is coded by the physical mask $\bm{M}\in\mathbb{R}^{W\times H}$, where $W$, $H$, and $N_{\lambda}$ denote the width, height, and the number of spectral channels, respectively. The coded HSI data cube is represented as ${\bf X}^{\prime}(:,:,n_{\lambda})={\bf X}(:,:,n_{\lambda})\odot{{\bf M}},n_{\lambda}=1,2,\ldots,N_{\lambda}$, where $\odot$ represents the element-wise multiplication. After light propagating through the disperser, each channel of ${\bf X}^{\prime}$ is shifted along the $\bm{H}$-${axis}$. The shifted data cube is denoted as ${\bf X}^{\prime\prime}\in\mathbb{R}^{W\times\tilde{H}\times N_{\lambda}}$, where $\tilde{H}=H+d_{\lambda}$. Here, $d_{\lambda}$ represents the shifted distance of the $N_{\lambda}$-th wavelength. This process can be formulated as modulating the shifted version $\tilde{{\bf X}}\in\mathbb{R}^{W\times\tilde{H}\times N_{\lambda}}$ with a shifted mask $\tilde{{{\bf M}}}\in\mathbb{R}^{W\times\tilde{H}\times N_{\lambda}}$, where $\bm{\tilde{M}}(i,j,n_{\lambda})=\bm{M}(w,h+d_{\lambda})$. At last, the imaging sensor captures the shifted image into a 2D measurement ${{\bf Y}}$, calculated as
$\vspace{-3mm}{{\bf Y}}=\textstyle\sum_{n_{\lambda}=1}^{N_{\lambda}}{\tilde{{\bf X}}}(:,:,n_{\lambda})\odot{\tilde{{{\bf M}}}}(:,:,n_{\lambda})+{\bf B},\vspace{-3mm}$ (1)
where ${\bf B}\in\mathbb{R}^{W\times\tilde{H}}$ denotes the measurement noise. By vectorizing the spectral data cube and measurement, that is $\bm{x}=\mathtt{vec}(\tilde{{\bf X}})\in\mathbb{R}^{\mathit{W\tilde{H}N_{\lambda}}}$ and $\bm{y}=\mathtt{vec}({{{\bf Y}}})\in\mathbb{R}^{\mathit{W\tilde{H}}}$, this model can be formulated as
$\vspace{-1mm}\bm{y}=\bm{A}\bm{x}+\bm{b},\vspace{-1mm}$ (2)
@@ 3.2 The Unfolding GAP Framework
Eq. (2) can be typically solved by convex optimization through the following objective:
@@
where $\tau$ is a noise-balancing factor. The first term guarantees that the solution $\hat{\bm{x}}$ fits the measurement, and the second term $R(\bm{x})$ refers to the image regularization.
To solve the optimization problem, we employ GAP (Generalized Alternating Projection) as our optimization framework, which extends classical alternating projection to the case in which projections are performed between convex sets that undergo a systematic sequence of changes. It can be interrupted anytime to return a valid solution and resumed subsequently to improve the solution [30]. This property is very suitable for deep unfolding which has very limited ‘optimization iterations’ (stages in the deep unfolding). Specifically, we introduce an auxiliary parameter ${\bm{v}}$, Eq. (3) can be written as:
$(\hat{\bm{x}},\hat{\bm{v}})=\underset{\bm{x},\bm{v}}{\arg\min}\frac{1}{2}\|\bm{x}-\bm{v}\|_{2}^{2}+\lambda R(\bm{v}),\quad\text{s.t.}\quad\bm{y}=\bm{A}\bm{x}.$ (4)
@@
$\bm{v}^{(k+1)}=\mathcal{D}_{k+1}(\bm{x}^{(k+1)}),$ (6)
where $\mathcal{D}_{k}$ is the neural network denoiser of the $k-th$ stage, and $\lambda$ is the noise penalty factor considered in learned denoisers. This step tries to map $\bm{x}^{(k+1)}$ to the target signal domain. Considering the projection step Eq. (5), assisted by deep network, we can modify it as follows:
$\bm{x}^{k+1}=\bm{v}^{(k)}+\text{DSC}(\bm{A}^{\top}(\bm{A}\bm{A}^{\top})^{-1}(\bm{y}-\bm{A}\bm{v}^{(k)})),$ (7)
where $\text{DSC}(\cdot)$ denotes a set of depthwise separable convolution and GELU [17] operations. The overall unfolding framework is shown in Fig. 3(d), where mask $A$ and measurement $y$ are inputs of the network. According to the Eqs. (5) and  (6), the first stage outputs $\bm{v}^{1}$ can be obtained. The detailed process is shown in Fig. 3(b). Typically, this process could make use of the position-specific degradation information in the projection part and close the gap between the sensing matrix and degradation matrix [28, 12]. Considering the stage number is much less than the iteration numbers in traditional model-based methods, it is difficult to achieve convergence with limited steps of gradient descent. However, these learnable linear and non-linear variations can be used to correct the gradients in these limited stage unfolding projections, and thus we name it Gradient Correlation GAP (GC-GAP).

### v2

@@ Abstract
Snapshot compressive spectral imaging reconstruction aims to reconstruct three-dimensional spatial-spectral images from a single-shot two-dimensional compressed measurement. Existing state-of-the-art methods are mostly based on deep unfolding structures but have intrinsic performance bottlenecks: $i$) the ill-posed problem of dealing with heavily degraded measurement, and $ii$) the regression loss-based reconstruction models being prone to recover images with few details. In this paper, we introduce a generative model, namely the latent diffusion model (LDM), to generate degradation-free prior to enhance the regression-based deep unfolding method by a two-stage training procedure. Furthermore, we propose a Trident Transformer (TT), which extracts correlations among prior knowledge, spatial, and spectral features, to integrate knowledge priors in deep unfolding denoiser, and guide the reconstruction for compensating high-quality spectral signal details. To our knowledge, this is the first approach to integrate physics-driven deep unfolding with generative LDM in the context of CASSI reconstruction. Comparisons between synthetic and real-world datasets illustrate the superiority of our proposed method in both reconstruction quality and computational efficiency. The code is available at https://github.com/Zongliang-Wu/LADE-DUN.
Keywords:
Spectral imaging Deep unfolding Diffusion model
{NoHyper}*Z. Wu and R. Lu—Contribute equally
@@ 1 Introduction
In contrast to normal RGB images which only have three spectral bands, hyperspectral images (HSIs) contain multiple spectral bands with more diverse spectral information. The spectral information serves to characterize distinct objects assisting high-level image tasks [60, 59, 35, 31, 50] and the observation of the world like medical imaging [38, 51] and remote sensing [18, 39]. However, the capture of HSIs is a question that has been studied for a long time because we need to collect 3-dimensional (3D) HSI signals by 2D sensors.
@@
Figure 1: (a) Comparison of PSNR (dB)-FLOPs (G) with previous HSI reconstruction methods. (b) The ablation study of using different time steps in diffusion. Our method achieves the desired results by only very few steps.
Figure 2: The top row: the error maps of the previous SOTA and our method. The bottom row is the feature map before and after applying LDM enhancement. The enhanced features demonstrate less noise and clearer edges.
@@
For many years, scientists have focused on how to collect HSIs in a quick and convenient method. In 2007, based on compressive sensing theory, a single-shot compressive spectral imaging way [17] was created to efficiently collect HSIs, named coded aperture snapshot spectral imaging (CASSI). The later improvement works [61, 44] provide better imaging quality and lower cost. CASSI modulates the HSI signal across various bands and combine all the modulated spectra to produce a 2D compressed measurement. Consequently, the task of reconstructing the 3D HSI signals from the 2D compressive measurements presents a fundamental challenge for the CASSI system.
@@
The reconstruction process can be viewed as solving an ill-posed problem. Many attempts at solving this problem including traditional model-based methods [2, 1, 70] and the learning-based methods[9, 47, 45] have been proposed since the inception of CASSI system. The deep unfolding network (DUN) is a combination of convex optimization and neural network prior (denoiser), enjoying both the interpretability of the model-based method and the power of learning-based methods. This branch of methods leads the development trend in recent years [64, 6, 46, 15, 32] and achieves SOTA performance.
@@
However, unlike super-resolution or deblurring that recovers from natural images, CASSI reconstruction has to recover HSIs from the compressed domain measurements, which results in severe degradation according to physical modulation, spectral compression, and multiple types of system noises. Thus, the CASSI reconstruction problem is much harder to learn intrinsic HSI properties than the normal image restoration tasks [66, 75, 40, 76, 33, 34, 30]. In the unfolding framework of the CASSI reconstruction method, the denoising network plays a critical role in deciding the final performance, which is embedded in each stage of the DUN. However, it always suffers from the performance bottleneck due to the intrinsic ill-posed problem of dealing with heavily degraded measurements. Thus, a high-performance denoiser with degradation-free knowledge is desired for CASSI reconstruction. Another problem is that previous popular regression-based reconstruction methods have difficulty in recovering details because the widely used regression losses are conservative with high-frequency details [55].
@@
To address these challenges, we introduce a generative prior in this paper to guide the reconstruction process in an unfolding framework. During training, the prior will be first learned from clean HSIs by an image encoder and then generated by a Latent Diffusion Model (LDM) from Gaussian noise and compressed measurement. Then, the learned prior is embedded into the deep denoiser of the DUN by a prior-guided Transformer. Significantly, our DUN is able to leverage external prior knowledge from clean HSIs and the powerful generative ability of LDM enhancing its reconstruction performance. The primary contributions presented in this paper can be summarized as follows:
@@
    We propose a novel LDM-based unfolding network for CASSI reconstruction, where the clean image priors are generated by a latent diffusion model to facilitate high-quality hyperspectral reconstruction. No additional data or training time is required. To the best of our knowledge, this is the first attempt to combine the physics-driven deep unfolding with generative LDM in CASSI reconstruction.
@@
    We design a three-in-one Transformer structure dubbed Trident Transformer (TT) to extract the correlation among prior knowledge, spatial, and spectral features. In TT, motivated by pansharpening techniques, we introduce an asymmetric cross-scale multi-head self-attention (ACS-MHSA) mechanism designed to efficiently fuse spatial-spectral features.
@@ 2.1 Diffusion Model in Low-level Vision
Diffusion models (DMs) [22, 56] are probabilistic generative models, which model the data distribution by learning a gradual iterative denoising process from the Gaussian distribution to the data distribution. Notably, they demonstrate promising capabilities in generating high-quality samples that encompass a wide range of modes, including super-resolution [16] and inpainting [41]. In light of the impressive achievements of diffusion models in image domains, numerous research endeavors [19, 21, 3, 23] have extended it to video generation. However, diffusion models suffer from significant computation inefficiency regarding data sampling, primarily due to the iterative denoising process required for inference. To address this challenge, several methods propose effective sampling techniques from trained diffusion models [57, 73], or alternatively learning the data distribution from a low-dimensional latent space [53], i.e. the latent diffusion model. The latent diffusion has a relatively faster speed and powerful generative ability for super-resolution and inpainting, but similar to the normal diffusion model, it is also prone to issues such as misaligned distribution of fine details and the occurrence of unwanted artifacts, leading to suboptimal performance in distortion-based metrics, e.g., PSNR. Moreover, latent diffusion costs large computational resources both for training and inference due to its large-size encoder and denoiser. Towards this end, some works combine the generative diffusion model with the regression restoration network and work well on distortion-based metrics like deblurring [52]. The recent works [12, 68] employ LDM on many low-level vision tasks and achieve SOTA with reasonable computational cost. We name these methods ‘integrated diffusion’ to distinguish them from the pure diffusion method. Nevertheless, employing diffusion models for the efficient reconstruction of hyperspectral images from highly compressed and degraded measurements presents significant challenges.
@@ 2.2 Hyperspectral Image Reconstruction
Before the advent of the deep learning wave, traditional model-based methods iteratively solved this inverse problem by convex optimization [63, 74, 62] with some hand-crafted constraints based on image priors, like sparsity [29] and low-rank [37]. These methods are robust and interpretable but require manual parameter tuning with low reconstruction speed and performance. With the help of deep learning, Plug-and-play (PnP) algorithms [8, 54, 71, 72, 49, 10, 11], embeds pre-trained denoising networks into convex optimization to solve the reconstruction problem, but still has limitations on performance because of the pre-trained denoiser. In recent years, the End-to-end (E2E) reconstruction directly trains a powerful deep neural network, like convolutional neural network (CNN) [13, 39, 25] and Transformers [5, 4, 7], to learn the recovery process from inputs (measurements) to outputs (desired HSIs). However, this simple design lacks interpretability and robustness for various hardware systems. Therefore, an interpretable design of a reconstruction network that unfolds a convex optimization process named DUN is proposed to leverage these problems. A series of CASSI reconstruction works based on DUN [43, 42, 6, 15, 69, 32] are proposed and become the state-of-the-art (SOTA) method. DUN can combine both interpretability in model-based methods and performance in deep learning-based methods to reconstruct CASSI at a fast speed. It changes iterative steps in optimization into several stages in a single network. The prior for optimization becomes a deep neural network denoiser. Since the DUN needs to define the forward model of imaging, it is also considered a physics-driven network. However, the recent DUNs still have bottlenecks for their regression-based denoiser design and the difficulty of dealing with compressive measurement features. Bearing these concerns, we propose an ‘integrated diffusion’ module and integrate it into the physics-driven DUN framework and design an efficient way to aggregate complex features during reconstruction.
@@
Figure 3: (a) The single disperser CASSI imaging process. HSI data cube is captured by a monochromatic sensor. (b) GC-GAP projection. (c) Latent encoder. (d) Simplified Denoiser. (e) The measurement $\bm{y}$ and masks $\bm{A}$ pass through an N-stage DUN, where each stage is composed of a GC-GAP projection and a denoiser. The denoiser follows a U-shape structure and consists of five Trident Transformers (TT), where each TT is assisted with prior knowledge $\bm{z}_{GT}$ generated from the diffusion model.
3 Problem Formulation
The CASSI system has high efficiency in capturing 3D spectral signals by initially coding spectral data with different wavelengths in an aperture and then integrating them into a 2D monochromatic sensor. The mathematical forward process of the widely used single-disperser CASSI (SD-CASSI) [62] can be illustrated as Fig. 3 (a). As can be seen, the original HSI data, denoted as $\bm{X}\in\mathbb{R}^{W\times H\times N_{\lambda}}$, is coded by the physical mask $\bm{M}\in\mathbb{R}^{W\times H}$, where $W$, $H$, and $N_{\lambda}$ denote the width, height, and the number of spectral channels, respectively. The coded HSI data cube is represented as ${\bf X}^{\prime}(:,:,n_{\lambda})={\bf X}(:,:,n_{\lambda})\odot{{\bf M}},n_{\lambda}=1,2,\ldots,N_{\lambda}$, where $\odot$ represents the element-wise multiplication. After light propagating through the disperser, each channel of ${\bf X}^{\prime}$ is shifted along the $\bm{H}$-${axis}$. The shifted data cube is denoted as ${\bf X}^{\prime\prime}\in\mathbb{R}^{W\times\tilde{H}\times N_{\lambda}}$, where $\tilde{H}=H+d_{\lambda}$. $d_{\lambda}$ is the shifted distance of the $N_{\lambda}$-th wavelength. This process can be formulated as modulating the shifted version $\tilde{{\bf X}}\in\mathbb{R}^{W\times\tilde{H}\times N_{\lambda}}$ with a shifted mask $\tilde{{{\bf M}}}\in\mathbb{R}^{W\times\tilde{H}\times N_{\lambda}}$, where $\bm{\tilde{M}}(i,j,n_{\lambda})=\bm{M}(w,h+d_{\lambda})$. At last, the imaging sensor captures the shifted image into a 2D measurement ${{\bf Y}}$, calculated as ${{\bf Y}}=\textstyle\sum_{n_{\lambda}=1}^{N_{\lambda}}{\tilde{{\bf X}}}(:,:,n_{\lambda})\odot{\tilde{{{\bf M}}}}(:,:,n_{\lambda})+{\bf B},$ where ${\bf B}\in\mathbb{R}^{W\times\tilde{H}}$ denotes the measurement noise. By vectorizing the data cube and measurement, that is $\bm{x}=\mathtt{vec}(\tilde{{\bf X}})\in\mathbb{R}^{\mathit{W\tilde{H}N_{\lambda}}}$ and $\bm{y}=\mathtt{vec}({{{\bf Y}}})\in\mathbb{R}^{\mathit{W\tilde{H}}}$, this model can be formulated as
$\bm{y}=\bm{A}\bm{x}+\bm{b},$ (1)
@@ 4 Proposed Model
To solve the problem in Eq. (1), we proposed a novel unfolding enhanced by latent diffusion prior. As shown in Fig. 3(e), in the inference phase, the measurement and masks pass through an N-stage DUN, where each stage is composed of a GC-GAP projection and a denoiser. The denoiser follows a U-shape structure and consists of Trident Transformers (Fig. 3(d)), where each TT is assisted with the LDM prior. We’ll describe these modules in more detail in this section.
@@ 4.1 The Unfolding GAP Framework
Eq. (1) can be typically solved by convex optimization by the objective below:
@@
where $\tau$ is a noise-balancing factor. The first term guarantees that the solution $\hat{\bm{x}}$ fits the measurement, and the term $R(\bm{x})$ refers to the image regularization.
To solve the optimization problem, we employ GAP (Generalized Alternating Projection) as our optimization framework, which extends classical alternating projection to the case in which projections are performed between convex sets that undergo a systematic sequence of changes. It can be interrupted anytime to return a valid solution and resumed subsequently to improve the solution [36]. This property is very suitable for DUN which has very limited ‘optimization iterations’ (stages in the DUN). Specifically, we introduce an auxiliary parameter ${\bm{v}}$, Eq. (2) can be written as:
$(\hat{\bm{x}},\hat{\bm{v}})=\underset{\bm{x},\bm{v}}{\arg\min}\frac{1}{2}\|\bm{x}-\bm{v}\|_{2}^{2}+\tau R(\bm{v}),\quad\text{s.t.}\quad\bm{y}=\bm{A}\bm{x}.$ (3)
@@
$\bm{v}^{(k+1)}=\mathcal{D}_{k+1}(\bm{x}^{(k+1)},\bm{z}),$ (5)
where $\mathcal{D}_{k}$ is the neural network denoiser of the $k-th$ stage and $\bm{z}$ is prior knowledge which will be described in Sec. 4.2. This step tries to map $\bm{x}^{(k+1)}$ to the target signal domain. Considering the projection step Eq. (4), assisted by deep network, we can modify it as follows:
$\bm{x}^{(k+1)}=\bm{v}^{(k)}+\text{DSC}(\bm{A}^{\top}(\bm{A}\bm{A}^{\top})^{-1}(\bm{y}-\bm{A}\bm{v}^{(k)})),$ (6)
where $\text{DSC}(\cdot)$ denotes a set of depthwise separable convolution and GELU [20] operations. The detailed process is shown in Fig. 3(b). Considering the stage number is much less than the iteration numbers in traditional model-based methods, it is difficult to achieve convergence with limited steps of gradient descent. Thus, we utilize these learnable parameters to rectify the gradients in the limited stage, and we refer to this method as the Gradient Correction GAP (GC-GAP). The overall unfolding framework is shown in Fig. 3(e), where mask $A$ and measurement $y$ are inputs of the network. According to the Eq. (6) and  (5), the first stage outputs $\bm{v}^{1}$ can be obtained.